6.4.25 Question Help - 10 -4 5 8 - 10 3 6 0 2 2 8 5 12 - 6 - 12 30 , is 3 -3 4 o . Find the QR factorization of A An orthogon#15
Answer #1

Solution:

Let

\small v_{_{1}}=\begin{bmatrix} -10\\ 2\\ -6\\ 16\\ 2\\ \end{bmatrix},\: \: v_{2}=\begin{bmatrix} 3\\ 3\\ -3\\ 0\\ 3\\ \end{bmatrix},\: \: v_{3}=\begin{bmatrix} 6\\ 0\\ 6\\ 6\\ 0\\ \end{bmatrix},\: \: v_{4}=\begin{bmatrix} 0\\ 4\\ 0\\ 0\\ -4\\ \end{bmatrix}

Now, normalize  \small v_{1},\: v_{2},\: v_{3},\: v_{4} to obtain an orthonormal basis:

\small u_{1}=\frac{v_{1}}{\left \| v_{1} \right \|}=\frac{v_{1}}{\sqrt{\left ( -10 \right )^{2}+2^{2}+\left ( -6 \right )^{2}+16^{2}+2^{2}}}=\frac{v_{1}}{\sqrt{400}}=\begin{bmatrix} -1/2\\\\ 1/10\\ \\ -3/10\\ \\ 4/5\\\\ 1/10 \end{bmatrix}

\small u_{2}=\frac{v_{2}}{\left \| v_{2} \right \|}=\frac{v_{2}}{\sqrt{3^{2}+3^{2}+\left ( -3 \right )^{2}+0^{2}+3^{2}}}=\frac{v_{2}}{6}=\begin{bmatrix} 1/2\\\\ 1/2\\ \\ -1/2\\ \\ 0\\\\ 1/2 \end{bmatrix}

\small u_{3}=\frac{v_{3}}{\left \| v_{3} \right \|}=\frac{v_{3}}{\sqrt{6^{2}+0^{2}+6^{2}+6^{2}+0^{2}}}=\frac{v_{3}}{6\sqrt{3}}=\begin{bmatrix} 1/\sqrt{3}\\\\ 0\\ \\ 1/\sqrt{3}\\ \\ 1/\sqrt{3}\\\\ 0 \end{bmatrix}

\small u_{4}=\frac{v_{4}}{\left \| v_{4} \right \|}=\frac{v_{4}}{\sqrt{0^{2}+4^{2}+0^{2}+0^{2}+\left ( -4 \right )^{2}}}=\frac{v_{3}}{4\sqrt{2}}=\begin{bmatrix} 0\\\\ 1/\sqrt{2}\\ \\0 \\ \\ 0\\\\ -1/\sqrt{2} \end{bmatrix}

\small \therefore Q=\begin{bmatrix} -1/2 & & 1/2 & 1/\sqrt{3} &0 \\\\ 1/10 & & 1/2 &0 &1/\sqrt{2} \\\\ -3/10 & & -1/2 &1/\sqrt{3} &0 \\\\ 4/5 & & 0 & 1/\sqrt{3} &0 \\\\ 1/10 & & 1/2 & 0& -1/\sqrt{2} \end{bmatrix}

\small R=Q^{T}A=\begin{bmatrix} -1/2 & & 1/2 & 1/\sqrt{3} &0 \\\\ 1/10 & & 1/2 &0 &1/\sqrt{2} \\\\ -3/10 & & -1/2 &1/\sqrt{3} &0 \\\\ 4/5 & & 0 & 1/\sqrt{3} &0 \\\\ 1/10 & & 1/2 & 0& -1/\sqrt{2} \end{bmatrix}^{T}\begin{bmatrix} -10 & -4 & 5 & 8 \\ 2 & 8 & 5 & 12 \\ -6 & -12& 3& 0 \\ 16 & 16 & 28 & 28 \\ 2 & 8 & 5 & 4 \end{bmatrix}

\small =\begin{bmatrix} 20 & 20 & 20 &20 \\ 0 & 12 &6 & 12\\ 0& 0 &12\sqrt{3} & 12\sqrt{3}\\ 0 & 0 &0 & 4\sqrt{2} \end{bmatrix}

The \small QR   factorization of  \small A is  \small A=QR , where

\small Q=\begin{bmatrix} -1/2 & & 1/2 & 1/\sqrt{3} &0 \\\\ 1/10 & & 1/2 &0 &1/\sqrt{2} \\\\ -3/10 & & -1/2 &1/\sqrt{3} &0 \\\\ 4/5 & & 0 & 1/\sqrt{3} &0 \\\\ 1/10 & & 1/2 & 0& -1/\sqrt{2} \end{bmatrix}

\small R=\begin{bmatrix} 20 & 20 & 20 &20 \\ 0 & 12 &6 & 12\\ 0& 0 &12\sqrt{3} & 12\sqrt{3}\\ 0 & 0 &0 & 4\sqrt{2} \end{bmatrix}